OpenStudy (liizzyliizz):

what is the general solution of (x^2-3) dy/dx = xy

OpenStudy (dumbcow):

\[y = k \sqrt{x^{2}-3}\]

OpenStudy (anonymous):

it can be written as 1/y dy=2x/2(x^2-3) dx integrate on both side ln(y)=lnsqrt(x^2-3)+lnc y=Csqrt|(x^2-3)